OpenStudy (anonymous):

On the Moon, a falling object falls just 2.65 feet in the first second after being dropped. Each second it falls 5.3 feet farther than in the previous second. How far would an object fall in the first ten seconds after being dropped?

OpenStudy (anonymous):

\[1.2.65+5.3=7.95\]\[2.7.95 + 5.3 =13.25\]can you keep going?

OpenStudy (anonymous):

3. ? + 5.3 = ? 4. ? + 5.3 = ? 5. ? + 5.3 = ? 6. ? + 5.3 = ? 7. ? + 5.3 = ? 8. ? + 5.3 = ? 9. ? + 5.3 = ? 10. ? + 5.3 = ?

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no its not finish

OpenStudy (anonymous):

please just tell me whats 3. 4. 5. 6. 7. 8. 9. 10. is? then we will keep going

OpenStudy (anonymous):

i will come back when you solve 3. 4. 5. 6. 7. 8. 9. 10. first ;)

OpenStudy (anonymous):

4. 18.55+5.3=24.05 so on and so on

OpenStudy (anonymous):

Thanks ; D

OpenStudy (anonymous):

yes and you need add all together 7.95 + 13.25 + ? + ? +? +? +? +? +?+? = ?

OpenStudy (anonymous):

and you Done! :)

OpenStudy (anonymous):

1. 2.65 2. 7.95 3. 13.25 4. 18.55 5. 23.85 6. 29.15 7. 34.45 8. 39.75 9. 45.05 10. 50.35 now add all together ;)

OpenStudy (anonymous):

i will come back to you when you finish and tell me the answer ;) and then you are done!